# Can Large Language Models Really Recognize Your Name?

### v1

@@ D.2 Human Evaluation
D.3 PII Detection Results
Name Type Predicted Type o3-mini DeepSeek R1 Claude 3.7 Sonnet DeepSeek V3 GPT-4o Gemini 1.5 Pro Gemini 2.0 Flash Qwen 2.5 7B Llama 3.1 8B Gemma 2 9B Flair PrivateAI
Location Person 95.55 97.66 97.28 97.87 85.09 85.99 97.54 80.27 94.81 96.73 92.89 98.6
Location 0.58 1.53 1.91 0.93 8.14 0.04 1.92 0.39 1.27 0.8 5.12 1.64
Org 0.02 0.56 0.0 0.08 0.27 0.0 0.01 9.4 0.5 0.0 0.1 0.96
Health 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 N/A N/A
Other 0.0 0.16 0.09 0.07 0.18 0.11 0.34 4.16 0.56 0.0 1.88 0.19
Missed 3.85 0.08 0.72 1.06 6.31 13.87 0.24 5.8 3.22 2.47 0.01 0.21
Org. Person 70.22 45.53 60.42 33.86 66.79 47.08 67.79 49.08 79.93 75.11 83.31 64.55
Location 0.01 0.01 0.01 0.0 0.12 0.07 0.01 0.0 0.0 0.01 0.01 1.9
Org 18.17 54.17 39.02 65.62 32.96 18.32 32.57 50.08 8.5 2.73 16.66 63.17
Health 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 N/A N/A
Other 0.01 0.16 0.1 0.06 0.04 1.38 0.11 0.04 0.39 0.0 0.0 0.13
Missed 11.6 0.16 0.45 0.46 0.09 33.15 0.15 0.87 11.7 22.22 0.03 0.56
Syndrome Person 96.8 95.65 81.08 87.21 74.4 65.4 89.6 85.7 97.35 89.82 84.37 73.05
Location 0.0 0.03 0.04 0.03 0.16 0.01 0.16 0.14 0.17 0.0 1.87 1.01
Org 0.03 0.42 0.25 0.8 0.25 0.02 0.7 6.12 0.17 0.16 1.61 8.63
Health 0.0 2.23 0.93 1.97 6.69 0.01 6.67 0.45 0.64 0.07 N/A N/A
Other 0.0 0.0 0.0 0.0 0.0 0.0 0.01 0.39 0.19 0.0 11.99 3.46
Missed 3.16 1.61 17.7 9.99 18.5 34.56 2.9 7.19 1.49 9.94 0.12 17.38
Mineral Person 55.41 60.9 20.52 14.47 10.19 3.2 39.2 41.95 60.1 19.44 40.63 18.33
Location 0.0 0.24 0.35 0.03 1.29 0.0 1.46 0.0 0.52 0.0 1.01 0.24
Org 0.03 29.57 0.63 0.28 3.97 0.0 7.3 11.83 0.42 0.03 0.21 1.08
Health 0.0 0.0 0.0 0.0 0.0 0.0 0.17 0.07 0.0 0.0 N/A N/A
Other 0.03 1.11 0.24 0.17 0.63 0.07 2.57 1.32 0.56 0.0 55.23 7.51
Missed 44.52 8.17 78.26 85.04 83.93 96.73 49.29 44.83 38.47 80.52 2.92 73.04
Bacterium Person 82.54 90.53 78.73 59.44 49.35 36.21 78.82 62.78 61.8 59.26 80.8 60.71
Location 0.0 0.36 0.95 0.0 3.49 0.03 2.75 0.24 0.89 0.0 2.69 0.92
Org 0.03 7.87 0.38 0.21 0.38 0.0 5.33 8.79 0.44 0.03 1.92 0.53
Health 0.0 0.0 0.0 0.0 0.0 0.0 0.03 0.3 0.09 0.0 N/A N/A
Other 0.03 0.12 0.03 0.06 0.38 0.24 2.57 2.22 15.36 0.0 14.56 11.63
Missed 17.4 1.12 19.91 40.3 46.39 63.52 10.5 25.71 21.42 40.71 0.03 26.63
Baseline Person 99.22 99.64 97.08 96.16 98.18 96.22 99.48 99.24 86.54 97.34 96.52 99.5
Location 0.12 0.28 0.6 0.52 0.94 0.36 0.56 0.1 0.44 0.54 1.74 0.7
Org 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.06 0.04 0.02 0.18 0.08
Health 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 N/A N/A
Other 0.0 0.04 0.0 0.14 0.0 0.0 0.0 0.04 3.48 0.02 1.56 0.06
Missed 0.66 0.04 2.32 3.18 0.88 3.42 0.0 0.56 9.5 2.1 0.0 0.06
@@ D.3 PII Detection Results
Table 6: Percentage of predicted entity types by each method for the different ambiguous name types and baseline. The “Other” category refers to predictions where the names are correct but the types are not included in Person, Location, Organization, or Health Info. “Missed” indicates that the names are not even detected (a small proportion of these cases is due to invalid output format). Note that Flair and PrivateAI do not have a “Health Info” category and can output multiple categories with different probabilities, causing the sum of their numbers to be greater than 1 in some cases.
Appendix E Additional Discussions
E.1 Clio
There are three main steps in Clio’s data preprocessing stage: 1) generating privacy-preserving summaries for (randomly sampled) Claude conversations, 2) clustering the summaries and generating cluster-level descriptions, and 3) filtering out clusters based on their sizes and the privacy auditor’s ratings. While it is shown empirically in Clio’s paper (Tamkin et al., 2024) that steps 2 and 3 can filter out the majority of leaked PIIs from step 1, we show one potential vulnerability by providing a sketch of a hypothetical attack that could be enabled by our paper’s methods:
  • Assume an attacker who has access to Clio’s final outputs (e.g., an employee) and wants to find more information about a Claude user with an ambiguous name (known to the attacker). Assume the user’s conversation is included in Clio’s inputs and their name is accidentally leaked in step 1 due to NRB and BPI.
  • The attacker, via Claude’s interface or API, creates a large volume of Claude conversations containing the target user’s name and “benign” prompt injection. This is meant to get the user’s conversation clustered in the same group as the attacker’s conversations (should have at least the minimum allowable cluster size).
  • The attacker searches in Clio’s final outputs for any cluster whose description contains the user’s name.
With this attack, a malicious actor can learn about the conversation topics of a particular user whose name happens to be confusable with another non-sensitive entity. If we cannot assume that the target user is included in Clio, then this attack can be used as a form of membership inference attack to check if the target is in Clio. This example highlights how LLMs’ inconsistent PII detection ability can open up novel attack avenues in systems that rely on this technology.

### v2

@@ D.2. Human Evaluation
D.3. Error Analysis
From Table 8, we can observe that the LLMs have an increased risk of misclassifying ambiguous human names as the same ambiguity source types. Specifically, location-like names have an increased risk of being classified as locations, organization-like as organization, syndrome-like as health, etc. Notably, in the case of minerals, bacteria, and syndrome, the LLMs have a very high rate of not detecting the name at all.
@@ D.3. Error Analysis
Table 8. Percentage of predicted entity types by each method for the different ambiguous name types and baseline. The ‘Other’ category refers to predictions where the names are correct but the types are not included in Person, Location, Organization, or Health Info. ‘Missed’ indicates that the names are not even detected (a small proportion of these cases is due to invalid output formatting). Note that Flair and PrivateAI do not have a ‘Health Info’ category and can output multiple categories with different probabilities, causing the sum of their numbers to be greater than 1 in some cases.
Name Type Pred. Type
DeepSeek R1
DeepSeek V3
GPT-4o
GPT-5
GPT-5 mini
Gemini 1.5 Pro
Gemini 2.5 Pro
Gemini 2.5 Flash
gpt-oss 20B
Qwen 2.5 7B
Llama 3.1 8B
Gemma 2 9B
Flair
PrivateAI
Location Per. 97.66 97.87 85.09 86.59 98.25 85.99 97.22 93.95 96.21 80.27 94.81 96.73 92.89 98.60
Loc. 1.53 0.93 8.14 0.77 1.22 0.04 1.88 3.52 1.53 0.39 1.27 0.80 5.12 1.64
Org. 0.56 0.08 0.27 0.07 0.05 0.00 0.11 1.13 0.21 9.40 0.50 0.00 0.10 0.96
Health 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Other 0.16 0.07 0.18 0.00 0.09 0.11 0.46 0.16 0.27 4.16 0.56 0.00 1.88 0.19
Missed 0.08 1.06 6.31 12.58 0.39 13.87 0.44 1.25 1.79 5.80 3.22 2.47 0.01 0.21
Organization Per. 45.53 33.86 66.79 59.52 64.66 47.08 67.60 59.09 65.39 49.08 80.22 75.11 83.31 64.55
Loc. 0.01 0.00 0.12 0.01 0.01 0.07 0.00 0.03 0.09 0.00 0.00 0.01 0.01 1.90
Org. 54.17 65.62 32.96 25.26 34.75 18.32 32.82 37.43 27.77 50.08 8.54 2.73 16.66 63.17
Health 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Other 0.16 0.06 0.04 0.08 0.47 1.38 0.61 0.14 1.72 0.04 0.79 0.00 0.00 0.13
Missed 0.16 0.46 0.09 15.12 0.11 33.15 0.12 3.31 5.03 0.87 10.98 22.22 0.03 0.56
Syndrome Per. 95.65 87.21 74.40 82.40 96.69 65.40 86.54 89.55 95.61 85.70 97.35 89.82 84.37 73.05
Loc. 0.03 0.03 0.16 0.01 0.04 0.01 0.15 0.12 0.12 0.14 0.17 0.00 1.87 1.01
Org. 0.42 0.80 0.25 0.00 0.35 0.02 0.53 0.67 0.24 6.12 0.17 0.16 1.61 8.63
Health 2.23 1.97 6.69 0.00 0.91 0.01 8.37 3.58 0.24 0.45 0.64 0.07 0.00 0.00
Other 0.00 0.00 0.00 0.00 0.00 0.00 0.02 0.00 0.01 0.39 0.19 0.00 11.99 3.46
Missed 1.61 9.99 18.50 17.59 1.98 34.56 4.51 6.07 3.79 7.19 1.49 9.94 0.12 17.38
Mineral Per. 60.90 14.47 10.19 12.03 57.53 3.20 47.30 33.22 37.88 41.95 60.10 19.44 40.63 18.33
Loc. 0.24 0.03 1.29 0.03 0.00 0.00 2.09 0.49 0.87 0.00 0.52 0.00 1.01 0.24
Org. 29.57 0.28 3.97 0.00 4.03 0.00 11.17 0.97 2.50 11.83 0.42 0.03 0.21 1.08
Health 0.00 0.00 0.00 0.00 0.00 0.00 0.07 0.00 0.00 0.07 0.00 0.00 0.00 0.00
Other 1.11 0.17 0.63 0.07 0.83 0.07 3.10 0.24 2.57 1.32 0.56 0.00 55.23 7.51
Missed 8.17 85.04 83.93 87.86 37.60 96.73 36.31 65.08 56.17 44.83 38.47 80.52 2.92 73.04
Bacterium Per. 90.53 59.44 49.35 71.15 93.31 36.21 89.17 77.93 92.13 62.78 61.80 59.26 80.80 60.71
Loc. 0.36 0.00 3.49 0.00 0.15 0.03 0.83 0.98 0.77 0.24 0.89 0.00 2.69 0.92
Org. 7.87 0.21 0.38 0.00 0.80 0.00 2.37 0.68 0.71 8.79 0.44 0.03 1.92 0.53
Health 0.00 0.00 0.00 0.00 0.00 0.00 0.53 0.00 0.00 0.30 0.09 0.00 0.00 0.00
Other 0.12 0.06 0.38 0.03 1.30 0.24 1.51 0.47 0.95 2.22 15.36 0.00 14.56 11.63
Missed 1.12 40.30 46.39 28.82 4.44 63.52 5.59 19.94 5.44 25.71 21.42 40.71 0.03 26.63
Baseline Per. 99.64 96.16 98.18 97.38 99.60 96.22 99.34 98.68 99.10 99.24 86.62 97.34 96.52 99.50
Loc. 0.28 0.52 0.94 0.04 0.22 0.36 0.30 0.66 0.56 0.10 0.44 0.54 1.74 0.70
Org. 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.06 0.04 0.02 0.18 0.08
Health 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Other 0.04 0.14 0.00 0.00 0.02 0.00 0.06 0.06 0.00 0.04 3.48 0.02 1.56 0.06
Missed 0.04 3.18 0.88 2.58 0.16 3.42 0.34 0.62 0.34 0.56 9.42 2.10 0.00 0.06
D.4. Additional PII Types
To demonstrate the generalizability of our approach, we also experiment with real usernames on the Reddit social network. Starting with an existing Reddit username dataset (Morris, 2017), we filter for usernames that satisfy the following criteria:
This process results in 4,408 unique usernames. Examples include: too-old-for-this, its-tough-out-there, dont-cancel-my-score, panda-with-a-hug, show-me-the-honey, … For the baseline, we use the top 200 usernames with the most comments on Reddit (excluding bot accounts).
Here is the prompt we use for assessing whether a username can be a compound modifier:
Your task is to check each phrase in the comma-separated list below to determine if it could be used as a potential hyphenated compound modifier that can modify a noun.
The phrases do not need to be well-known compound modifiers as long as they can be used in a sentence, perhaps in a creative manner.
If a phrase contains a real human name, or is offensive or sexually suggestive, do not include it.
Here is the list of phrases: [<PHRASE_1>, <PHRASE_2>, ...]
Examples of phrases that can form a meaningful compound modifier:
loud-mouthed => He is a loud-mouthed person.
step-by-step => She has a step-by-step approach to problem solving.
state-of-the-art => They managed to develop a state-of-the-art solution.
First, try to construct a short sentence using the phrase as a compound word creatively. Then, reason about whether the meaning of the sentence is clear and interpretable by a human.
Finally, output valid phrases in JSON format at the end like this: {"output": [<PHRASE_1>, <PHRASE_2>, ...]}
Start here:
We use the following five templates:
As a veteran Reddit user, I cant believe I got to experience such a surreal [MASK] moment. The owner of that account must be very excited about the attention right now.
I have been using Reddit for a long time, and I finally managed to participate in a genuine [MASK] event. That account only invited its top 10 commenters.
I recently joined Reddit to see what its about, and I had to deal with this [MASK] person. The user was so annoying, I just want to block their account.
I just created an account on Reddit, and immediately got into this [MASK] situation. I honestly dont know what to do, such Redditors are beyond my understanding.
I just witnessed yet another [MASK] episode on my Reddit account. At this point, the user should charge for their high-quality drama.
We also evaluate how the examined mitigations can impact the username detection performance of three representative LLMs. Overall, the improvement is much lower compared to the human names.
Table 9. Average Recall (R) and False Discovery Rate (FDR) (formatted in percentage) of Reddit usernames for three representative LLMs with different “defenses”. Takeaway: Recall improves significantly with defenses, but remains below 0.65. FDR does not improve for gpt-oss-20B.
Name type Prompt Method GPT-5-mini Gemini 2.5 Flash gpt-oss-20B
[…]
Compound Orig. 0.37 0.52 0.20 39.21 0.194 3.71
P1 0.57 0.21 0.24 6.62 0.34 4.95
P2 0.65 0.09 0.29 5.74 0.36 7.03
Baseline Orig. 0.87 0.00 0.82 2.85 0.62 1.43
P1 0.90 0.00 0.87 1.37 0.71 0.84
P2 0.93 0.00 0.88 0.68 0.76 0.52